# Reentrant Superconductivity in Zeeman Fields

### v1

@@ Abstract
Introduction: An external magnetic field suppresses superconductivity in two ways. The orbital effect fluctuates the phase of superconducting condensate and increases the free-energy of a superconducting state. This effect is common for all superconductors (SCs). In layered SCs or monolayer SCs, the orbital effect is negligible for a magnetic field applied parallel to the layers. Only the Zeeman effect modifies the superconducting states. A Zeeman field always suppresses spin-singlet superconductivity. The critical magnetic field $H_{c}$ at zero temperature is called Pauli limit $H_{\mathrm{P}}$[14, 4]. For spin-triplet superconductivity, the effect depends on the alignment between the direction of a Zeeman filed $\bm{H}$ and that of a $\bm{d}$ vector. A Zeeman filed parallel to $\bm{d}$ suppresses superconductivity, whereas that perpendicular to $\bm{d}$ does not affect the superconducting state. In experiments, Zeeman-field-induced superconductivity (ZFIS) or reentrant superconductivity have been observed various materials such as an organic SC $\lambda-\mathrm{(BETS)}_{2}\mathrm{FeCl}_{4}$ [20], Europium compounds [10], UCoGe[2], and UTe2 [12, 1, 7]. The last one has attracted attention these days as a promising candidate of a spin-triplet SC. However, our understanding of ZFIS remains extremely limited. Only the Jaccarino-Peter compensation [8] is the widely accepted as the mechanism explaining the transition from a magnetic phase to a superconducting phase. Thus, another scenario for explaining the phenomenon is desired to understand physics behind the experimental findings. We will address such an issue in this paper.
@@
To demonstrate the ZFIS, we introduce the third vector in spin space $\bm{\alpha}$ that describes spin-orbit interactions (SOIs). A Zeeman field and a SOI induce additional pairing correlations. Although they are not linked to the pair potentials, they govern the stability of superconductivity depending on their frequency symmetry classes. Odd-frequency Cooper pairs decrease the transition temperature $T_{c}$ [3, 11, 19] because they decrease the superfluid weight. On the other hand, even-frequency Cooper pairs increase $T_{c}$[16] In this paper, we show that spin-singlet Cooper pairs whose correlation function is proportional to a scaler chiral product $\bm{\alpha}\times\bm{H}\cdot\bm{d}$ assist spin-triplet superconductivity in sufficiently strong Zeeman fields. We use the unit of $k_{B}=\hbar=c=1$, where $k_{B}$ is the Boltzmann constant and $c$ is the speed of light.
@@
Figure 1: The critical magnetic field $H_{c}$ is plotted as a function of temperature $T$. (a) All the three vectors are align in parallel, $\bm{d}\parallel\bm{\alpha}\parallel\bm{H}$. The arrow at the vertical axis indicates the Pauli limit $H_{\mathrm{P}}$. (b) The three vectors are perpendicular to one another, $\bm{d}\perp\bm{\alpha}$, $\bm{d}\perp\bm{H}$, and $\bm{\alpha}\perp\bm{H}$. At $H=0$, superconductivity is absent because of $\alpha>\mu_{\mathrm{B}}H_{\mathrm{P}}$. (c) The mixed states with $\alpha_{x}=0.75T_{0}$ exhibit the reentrant superconductivity. The two superconducting phases tend to separate from each other for large $\alpha_{y}$.
@@
Model: We consider a electronic structure in two dimension, where two Fermi surfaces surrond $\bm{K}$ and $\bm{K}$ in momentum space. We describe such two Fermi surfaces in terms of two valleys. The normal state Hamiltonian is given by
$\check{H}_{\mathrm{N}}$ $(\bm{k})=\hat{\xi}_{\bm{k}}\hat{\sigma}_{0}+\bm{\alpha}\cdot\hat{\bm{\sigma}}\hat{\rho}_{z}+\mu_{\mathrm{B}}\bm{H}\cdot\hat{\bm{\sigma}}\hat{\rho}_{0},$ (1)
$\hat{\xi}_{\bm{k}}$ $=\frac{1}{2m}\left(\bm{k}\hat{\rho}_{0}-\bm{K}\hat{\rho}_{z}\right)^{2}-\mu,$ (2)
@@
where $\mu$ is the chemical potential and $\bm{H}$ represents a Zeeman field with $\mu_{\mathrm{B}}$ being the Bohr magneton. The Pauli matrices in spin and valley spaces are denoted by $\hat{\bm{\sigma}}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$ and $\hat{\bm{\rho}}=(\hat{\rho}_{x},\hat{\rho}_{y},\hat{\rho}_{z})$, respectively. The unit matrices in the two spaces are denoted by $\hat{\sigma}_{0}$ and $\hat{\rho}_{0}$. We assume that the SOI $\bm{\alpha}$ is independent of $\bm{k}$ and changes its sign in the two valleys. The time-reversal operation is given by $\mathcal{T}=i\hat{\sigma}_{y}\hat{\rho}_{x}\mathcal{K}$, where $\mathcal{K}$ means the complex conjugation plus $\bm{k}\to-\bm{k}$. The particle-hole conjugation is represented by $\undertilde{\check{H}}_{\mathrm{N}}(\bm{k})=\check{H}^{*}_{\mathrm{N}}(-\bm{k})$ as usual. Two electrons at the different valleys form a spin-triplet Cooper pair. The pair potential for such a pair is represented as
@@
where $\bm{d}$ is independent of $\bm{k}$. Instead of $\bm{d}$ being odd-parity function, making valley-parity odd preserves the antisymmetric property of the pair potential derived from the Fermi-Dirac statistic of electrons. We note that all Cooper pairs in this model belong to even-parity $s$-wave symmetry class. Although the similar electronic structures are realized in transition-metal dicalcogenides [15, 21, 5], we do not focus on any specific materials in this paper. The advantage of the model is briefly summarized as follows. In single-band SCs, $\bm{\alpha}$ and $\bm{d}$ are an odd-parity function depends on $\bm{k}$. In our model, they are described by the $\bm{k}$ independent potentials which change sign under interchanging the valley indices. As a result, magnetic properties of a SC is purely derived from the relative vector configuration among $\bm{d}$, $\bm{\alpha}$, and $\bm{H}$. At the end of this paper, we will confirm that main conclusions of this paper are valid also in single-band odd-parity spin-triplet SCs.
@@
$\left[i\omega_{n}-H_{\mathrm{BdG}}(\bm{k})\right]\begin{bmatrix}\check{\mathcal{G}}&\check{\mathcal{F}}\\ -\undertilde{\check{\mathcal{F}}}&-\undertilde{\check{\mathcal{G}}}\end{bmatrix}_{(\bm{k},\omega_{n})}=1,$ (4)
@@
where $\omega_{n}=(2n+1)\pi T$ is a Matsubara frequency with $T$ being a temperature. As shown in Eq. (17) in Appendix A, the Hamiltonian is block-diagonalized into two particle-hole spaces. The analytical expression of the anomalous Green’s function is supplied in Eq. (19).
Table 1: Symmetry classification of the pair correlation functions. The top row corresponds to the spin-triplet pairing correlation that is linked to the pair potential through the gap equation. It appears at the first term in Eqs. (6) and (9). The second and third low represent odd-frequency pairing correlations appearing at the second term in Eqs. (6) and (9). They make the superconducting unstable. The cooperative effect between a Zeeman filed and a SOI induces the spin-singlet pairing correlation at the bottom row. All the Cooper pairs belong to even-momentum-parity $s$-wave symmetry class.
@@
even triplet $\bm{d}\cdot\hat{\bm{\sigma}}$ odd $\hat{\rho}_{y}$ pair potential
odd singlet $\bm{d}\cdot\bm{H}$ odd $\hat{\rho}_{y}$ $F_{\parallel}$
@@
odd triplet $\bm{\alpha}\times\bm{d}\cdot\hat{\bm{\sigma}}$ even $\hat{\rho}_{x}$ $F_{\perp}$
even singlet $\bm{\alpha}\times\bm{d}\cdot\bm{H}$ even $\hat{\rho}_{x}$ $F_{\perp}$
@@
Results: At first, we discuss a configuration $\bm{\alpha}\parallel\bm{H}\parallel\bm{d}$, where the three vectors align in the same direction. The anomalous Green’s function is given in Eqs (26). To solve the linearized gap equation, the summation over $\bm{k}$ is necessary. The shift of the wavenumber $\bm{k}\to\bm{k}\mp\bm{K}$ for $\xi_{\pm}$ remains the results unchanged because $\bm{\alpha}$ and $\bm{d}$ are independent of $\bm{k}$. By changing the summation over $\bm{k}$ to the integration over $\xi_{\bm{k}}\mp\alpha\hat{\rho}_{z}$, we reach an expression of
$\frac{1}{V_{\mathrm{vol}}}\sum_{\bm{k}}$ $\check{F}_{\parallel}(\bm{k},\omega_{n})=-\frac{N_{0}\pi}{|\omega_{n}|}\left[\frac{\omega_{n}^{2}}{\omega_{n}^{2}+\mu_{\mathrm{B}}^{2}H^{2}}\bm{d}\cdot\hat{\bm{\sigma}}\right$.
$\left.-\frac{i\,\omega_{n}\,\mu_{\mathrm{B}}\bm{d}\cdot\bm{H}}{\omega_{n}^{2}+\mu_{\mathrm{B}}^{2}H^{2}}\right]\left(i\hat{\sigma}_{y}\right)(i\hat{\rho}_{y}),$ (6)
@@
where $N_{0}$ is the density of states at the Fermi level per spin. The SOI modifies the band-dispersion only slightly in this configuration. The first term contributes to the pair potential through the gap equation in Eq. (23). A Zeeman field decreases its amplitude and generates odd-frequency pairing correlation appearing at the second term. Odd-frequency pairing correlations makes the superconducting state unstable [3]. The two pairing correlations are listed at the first and second rows in Table 1. We find that the structure of the anomalous Green’s function in Eq. (6) is essentially the same as that for a spin-singlet superconductor under a Zeeman field by changing $\bm{d}\cdot\hat{\bm{\sigma}}\to\Delta$ and $\bm{d}\cdot\bm{H}\to\Delta\bm{H}\cdot\hat{\bm{\sigma}}$. In fact, the resulting linearized gap equations in the two cases
$\ln\frac{T}{T_{0}}+2\pi T\sum_{\omega_{n}>0}\frac{1}{\omega_{n}}\left[1-\frac{\omega_{n}^{2}}{\omega_{n}^{2}+\mu_{\mathrm{B}}^{2}H^{2}}\right]=0,$ (7)
are identical to each other. Here $T_{0}$ is the transition temperature at $\bm{H}=\bm{\alpha}=0$. We draw the HTH-T phase diagram shown in Fig. 1(a) in two ways: solving the gap equation and finding the minimum of the free-energy,
$\Omega_{S}=$ $-\frac{1}{2}T\frac{1}{V_{\mathrm{vol}}}\sum_{\bm{k}}\sum_{j=1}^{4}\log\left(2\cosh\left[\frac{E_{\bm{k},j}}{2T}\right]\right)+\frac{\bm{d}^{2}}{g}$
$+\text{Const.},$ (8)
@@
where $E_{\bm{k},j}$ is the eigen energy of the BdG Hamiltonian. Two results are identical to each other when the transition to the superconducting states is a second-order. For first-order transitions, the phase boundary is obtained from the free-energy minima. The transition to the superconducting state becomes a first-order for $T<0.556T_{0}$ [9, 17] because odd-frequency pairing correlation decreases the coefficient of the fourth order term in the Ginzburg-Landau expansion and changes its sign to negative at low temperatures [18]. The arrow on the vertical axis indicate the Pauli limit $H_{\mathrm{P}}$ [14, 4]. A Zeeman field seriously suppresses spin-triplet superconductivity in the parallel configuration.
@@
Secondly, we consider a configuration $\bm{\alpha}\perp\bm{d}$, $\bm{\alpha}\perp\bm{H}$, and $\bm{H}\perp\bm{d}$, where the three vectors are perpendicular to one another. The anomalous Green’s function near the transition temperature is supplied in Appendix (28). Here we show the results after carrying out the summation over $\bm{k}$ by shifting the wavenumber $\bm{k}\to\bm{k}\mp\bm{K}$ for $\xi_{\pm}$,
$\frac{1}{V_{\mathrm{vol}}}\sum_{\bm{k}}$ $\check{F}_{\perp}(\bm{k},\omega_{n})=-\frac{N_{0}\pi}{|\omega_{n}|(\omega_{n}^{2}+\alpha^{2}+\mu_{\mathrm{B}}^{2}H^{2})}$
$\times\left[(\omega_{n}^{2}+\mu_{\mathrm{B}}^{2}H^{2})\bm{d}\cdot\hat{\bm{\sigma}}+\omega_{n}(\bm{\alpha}\times\bm{d})\cdot\hat{\bm{\sigma}}\hat{\rho}_{z}\right$.
$\left.+i(\bm{\alpha}\times\bm{d})\cdot\mu_{B}\bm{H}\hat{\rho}_{z}\right]i\hat{\sigma}_{y}(i)\hat{\rho}_{y}.$ (9)
@@
The first term is the pairing correlation linked to the pair potential. The odd-frequency pairing correlation at the second term implies that the SOI suppresses superconductivity [6]. The third term plays a key role in the ZFIS. This term stabilizes the superconducting state at high Zeeman field because it belongs to even-frequency symmetry class. The linearized gap equation results in
$\ln\frac{T}{T_{0}}$ $+2\pi T\sum_{\omega_{n}>0}\frac{1}{\omega_{n}}\left[1-\frac{\omega_{n}^{2}+\mu_{\mathrm{B}}^{2}H^{2}}{\omega_{n}^{2}+\mu_{\mathrm{B}}^{2}H^{2}+\alpha^{2}}\right]=0.$ (10)
@@
Comparing the results with Eq. (7) suggests that the large SOI $\alpha>\mu_{\mathrm{B}}H_{\mathrm{P}}$ deletes superconductivity at $\bm{H}=0$. In the HTH-T phase diagram in Fig. 1(b), superconductivity is absent at $\bm{H}=0$ because we choose such large SOI as $\alpha=1.5T_{0}$ and $2.5T_{0}$. Equation (10) also indicates that a Zeeman field screens such negative effects due to the SOI. Fig. 1(b) shows that the superconducting phase appears in large Zeeman fields and that the critical temperature increases with increasing Zeeman fields. The induced spin-singlet Cooper pairs at the last term in Eq. (9) stabilize the spin-triplet superconducting states at high Zeeman fields. We conclude that the interplay between the SOI and the Zeeman field realizes ZFIS. Eq. (10) suggests that $T_{c}$ goes to the $T_{0}$ for $\mu_{\mathrm{B}}H\gg\alpha$. However, our phenomenological theory cannot predict such limiting behavior. We have assumed in high magnetic fields that the attractive interaction between two electrons remains unchanged and that other conduction bands do not come to the Fermi level. In Table 1, we list the pairing correlations appearing in a SC and their pairing symmetry.
@@
Next, we consider mixed situations between the parallel and the perpendicular configurations,
$\bm{\alpha}=$ $\alpha_{x}\bm{e}_{x}+\alpha_{y}\bm{e}_{y},$ (11)
$\bm{d}=$ $d_{x}\bm{e}_{x}+d_{z}\bm{e}_{z},$ (12)

### v3

@@ Abstract
Introduction: An external magnetic field suppresses superconductivity in two ways: orbital effect and spin Zeeman effect. The orbital effect common to all superconductors (SCs) fluctuates the phase of the superconducting condensate and increases the free-energy. In layered SCs and monolayer SCs under a magnetic field applied parallel to the layers, the orbital effect is negligible and only the Zeeman effect modifies the superconducting states. A Zeeman field always suppresses spin-singlet superconductivity. The critical magnetic field $H_{c}$ at zero temperature is called Pauli limit $H_{\mathrm{P}}$[19, 6]. For spin-triplet superconductivity, a Zeeman filed $\bm{H}$ parallel to a $\bm{d}$ vector suppresses superconductivity, whereas that perpendicular to the $\bm{d}$ does not affect the superconducting state. In experiments, Zeeman-field-induced superconductivity (ZFIS) or reentrant superconductivity have been observed various materials such as an organic SC $\lambda-\mathrm{(BETS)}_{2}\mathrm{FeCl}_{4}$ [26], Europium compounds [15], UCoGe[2], LaTiO3-KTaO3 interface [14], and UTe2 [17, 1, 9]. The last one has attracted attention these days as a promising candidate of a spin-triplet SC. However, our understanding of ZFIS remains extremely limited. Only the Jaccarino-Peter compensation [10] is the widely accepted as the mechanism explaining the transition to a superconducting phase in high Zeeman fields. Thus, another scenario for explaining the ZFIS is desired to understand physics behind the experimental findings. We will address such an issue in this paper.
@@
To demonstrate the ZFIS, we introduce the third vector in spin space $\bm{\alpha}$ that describes spin-orbit interactions (SOIs). It has been established that SOIs increase $H_{c}$ of spin-singlet SCs beyond the Pauli limit [11, 24]. A recent paper reported very complicated phase diagrams for spin-triplet SCs with SOIs [12]. A Zeeman field and a SOI induce additional pairing correlations. Although they do not form the pair potential, they govern the stability of superconductivity depending on their frequency symmetry classes. Odd-frequency Cooper pairs decrease the transition temperature $T_{c}$ [5, 16, 25] because they decrease the superfluid weight. On the other hand, even-frequency Cooper pairs increase $T_{c}$[21] In this paper, we show that spin-singlet Cooper pairs whose correlation function is proportional to a scalar chiral product $\bm{\alpha}\times\bm{V}\cdot\bm{d}$ assist spin-triplet superconductivity in sufficiently strong Zeeman fields.
@@
Figure 1: The critical magnetic field $H_{c}$ is plotted as a function of temperature $T$. (a) All the three vectors are aligned in parallel, $\bm{d}\parallel\bm{\alpha}\parallel\bm{H}$. The arrow at the vertical axis indicates the Pauli limit $H_{\mathrm{P}}$. (b) The three vectors are perpendicular to one another, $\bm{d}\perp\bm{\alpha}$, $\bm{d}\perp\bm{H}$, and $\bm{\alpha}\perp\bm{H}$. At $H=0$, superconductivity is absent because of $\alpha>\mu_{\mathrm{B}}H_{\mathrm{P}}$. (c) The mixed states with $\alpha_{x}=0.75T_{0}$ exhibit the reentrant superconductivity. The two superconducting phases tend to separate from each other for large $\alpha_{y}$.
@@
Model: We consider an electronic structure in two dimension, where two Fermi surfaces surround at certain points $\bm{K}$ and $\bm{K}$ in momentum space. We describe such two Fermi surfaces in terms of two valleys. The normal state Hamiltonian is given by
$\check{H}_{\mathrm{N}}$ $(\bm{k})=\hat{\xi}_{\bm{k}}\hat{\sigma}_{0}+\bm{\alpha}\cdot\hat{\bm{\sigma}}\hat{\rho}_{z}+\bm{V}\cdot\hat{\bm{\sigma}}\hat{\rho}_{0},$ (1)
$\hat{\xi}_{\bm{k}}$ $=\frac{1}{2m}\left(\bm{k}\hat{\rho}_{0}-\bm{K}\hat{\rho}_{z}\right)^{2}-\mu,\quad\bm{V}=\mu_{\mathrm{B}}\bm{H}$ (2)
@@
where $\mu$ is the chemical potential and $\bm{H}$ represents a Zeeman field with $\mu_{\mathrm{B}}$ being the Bohr magneton. We use the unit of $k_{B}=\hbar=c=1$, where $k_{B}$ is the Boltzmann constant and $c$ is the speed of light. The Pauli matrices in spin and valley spaces are denoted by $\hat{\bm{\sigma}}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$ and $\hat{\bm{\rho}}=(\hat{\rho}_{x},\hat{\rho}_{y},\hat{\rho}_{z})$, respectively. The unit matrices in the two spaces are denoted by $\hat{\sigma}_{0}$ and $\hat{\rho}_{0}$. We assume that the SOI $\bm{\alpha}$ is independent of $\bm{k}$ and changes its sign in the two valleys. The time-reversal operation is given by $\mathcal{T}=i\hat{\sigma}_{y}\hat{\rho}_{x}\mathcal{K}$, where $\mathcal{K}$ means the complex conjugation plus $\bm{k}\to-\bm{k}$. The particle-hole conjugation is represented by $\undertilde{\check{H}}_{\mathrm{N}}(\bm{k})=\check{H}^{*}_{\mathrm{N}}(-\bm{k})$ as usual. Two electrons at the different valleys form a spin-triplet Cooper pair. The pair potential for such a pair is represented as
@@
where $\bm{d}$ is independent of $\bm{k}$. Instead of $\bm{d}$ being odd-parity function, making valley-parity odd preserves the antisymmetric property of the pair potential derived from the Fermi-Dirac statistics of electrons. We note that all Cooper pairs in this model belong to even-parity $s$-wave symmetry class. Although the similar electronic structures are realized in transition-metal dicalcogenides [20, 27, 7], we do not focus on any specific materials in this paper. The advantage of the model is briefly summarized as follows. In single-band SCs, $\bm{\alpha}$ and $\bm{d}$ are an odd-parity function depends on $\bm{k}$. In our model, they are described by the $\bm{k}$ independent potentials which change the sign under interchanging the valley indices. As a result, magnetic properties of a SC is purely derived from the relative vector configuration among $\bm{d}$, $\bm{\alpha}$, and $\bm{H}$. At the end of this paper, we will confirm that main conclusions of this paper are valid also in single-band odd-parity spin-triplet SCs.
@@
$\left[i\omega_{n}-H_{\mathrm{BdG}}(\bm{k})\right]\begin{bmatrix}\check{{G}}&\check{{F}}\\ -\undertilde{\check{{F}}}&-\undertilde{\check{{G}}}\end{bmatrix}_{(\bm{k},\omega_{n})}=1,$ (4)
@@
where $\omega_{n}=(2n+1)\pi T$ is a Matsubara frequency with $T$ being a temperature. The analytical expression of the anomalous Green’s function is supplied in Eq. (S4) in Supplemental Material [3]. By substituting the anomalous Green’s function into the gap equation in Eq. (S10), the pair potential is calculated in a self-consistent way.
The thermodynamics of a superconductor near the transition temperature is described well by the Ginzburg-Landau free-energy,
$\Omega_{S}=$ $-\frac{1}{2}T\frac{1}{V_{\mathrm{vol}}}\sum_{\bm{k}}\sum_{j=1}^{4}\log\left(2\cosh\left[\frac{E_{\bm{k},j}}{2T}\right]\right)+\frac{\bm{d}^{2}}{g},$ (6)
$\approx$ $\,a\,d^{2}+b\,d^{4}+\mathrm{h.o.t},$ (7)
where $E_{\bm{k},j}$ is the eigen energy of the BdG Hamiltonian and $d$ in Eq. (7) is the amplitude of the $\bm{d}$ vector. The coefficient $a$ is proportional to the linearized gap equation which is 0 at $T=T_{c}$ and negative for $T<T_{c}$ as shown below in Eqs. (11) and (13). Magnetically active potentials generate the pairing correlations belonging to various symmetry classes as shown in Eq. (S4). Among them, only the spin-triplet odd-valley parity pairing correlation form the pair potential through the gap equation. Hereafter, we refer to such pairing correlation as principal correlation. Although the decrease in $T_{c}$ certainly provides a quantitative measure of the instability of the superconducting state, the information obtained from the gap equation is limited to the amplitude of the principal correlation. To compensate for the lack of information, we discuss how remaining correlations in Eq. (S4) stabilize or destabilize the superconducting states by using the superfluid weight defined by
$Q_{F}=T\sum_{\omega_{n}}\int d\xi_{\bm{k}}\,\frac{1}{2}\mathrm{Tr}[-F\undertilde{F}]_{(\bm{k},\omega_{n})}.$ (8)
In SM [3], a relation of $Q_{F}$ to the Meissner screening length is also explained. We mainly discuss the superfluid weight at the lowest order of $\bm{d}$,
$q(T,H)=\left.\frac{Q_{F}(T,H)}{d^{2}}\right|_{d\to 0}.$ (9)
Here we summarize several general features of $q$. The superfluid weight of induced odd-frequency pairing correlations is negative. As a consequence, the appearance of odd-frequency pairs makes the superconducting states unstable and decreases $T_{c}$ [5]. Such situation is indirectly reflected in the coefficient $a$ in Eq. (7). The coefficient $b$ in Eq. (7) is proportional to $q(T_{c},H_{c})$ [23]. Therefore, the transition to the superconducting phase becomes a first-order for $q(T_{c},H_{c})<0$.
Table 1: Symmetry classification of the pair correlation functions. The top row corresponds to the spin-triplet pairing correlation that is linked to the pair potential through the gap equation and is referred to as principal pairing correlation. It appears at the first term in Eqs. (10) and (12). The second and third low represent odd-frequency pairing correlations appearing at the second term in Eqs. (10) and (12). They make the superconducting unstable. The cooperative effect between a Zeeman filed and a SOI induces the spin-singlet pairing correlation at the bottom row. All the Cooper pairs belong to even-momentum-parity $s$-wave symmetry class.
@@
even triplet $\bm{d}\cdot\hat{\bm{\sigma}}$ odd $\hat{\rho}_{y}$ principal
odd singlet $\bm{d}\cdot\bm{H}$ odd $\hat{\rho}_{y}$ induced $F_{\parallel}$
@@
odd triplet $\bm{\alpha}\times\bm{d}\cdot\hat{\bm{\sigma}}$ even $\hat{\rho}_{x}$ induced $F_{\perp}$
even singlet $\bm{\alpha}\times\bm{d}\cdot\bm{H}$ even $\hat{\rho}_{x}$ induced $F_{\perp}$
@@
Results: At first, we discuss a configuration $\bm{\alpha}\parallel\bm{H}\parallel\bm{d}$, where the three vectors align in the same direction. The anomalous Green’s function is given in Eq. (S13) in SM. To solve the linearized gap equation, the summation over $\bm{k}$ is necessary. The shift of the wavenumber $\bm{k}\to\bm{k}\mp\bm{K}$ for $\xi_{\pm}$ remains the results unchanged because $\bm{\alpha}$ and $\bm{d}$ are independent of $\bm{k}$. By changing the summation over $\bm{k}$ to the integration over $\xi_{\bm{k}}\mp\alpha\hat{\rho}_{z}$, we reach an expression of
$\frac{1}{V_{\mathrm{vol}}}$ $\sum_{\bm{k}}\check{F}_{\parallel}(\bm{k},\omega_{n})$
$=$ $\frac{N_{0}\pi}{|\omega_{n}|}\left[\frac{\omega_{n}^{2}}{\omega_{n}^{2}+V^{2}}\bm{d}\cdot\hat{\bm{\sigma}}-\frac{i\,\omega_{n}\,\bm{d}\cdot\bm{V}}{\omega_{n}^{2}+V^{2}}\right]\hat{\sigma}_{y}\hat{\rho}_{y},$ (10)
@@
where $N_{0}$ is the density of states at the Fermi level per spin. The SOI modifies the band-dispersion only slightly in this configuration. The first term corresponds to the principal correlation. A Zeeman field generates the odd-frequency pairing correlation at the second term. The two pairing correlations are listed at the first and second rows in Table 1. We find that the structure of the anomalous Green’s function in Eq. (10) is essentially the same as that for a spin-singlet superconductor under a Zeeman field by changing $\bm{d}\cdot\hat{\bm{\sigma}}\to\Delta$ and $\bm{d}\cdot\bm{V}\to\Delta\bm{V}\cdot\hat{\bm{\sigma}}$. In fact, the resulting linearized gap equations in the two cases
$\ln\frac{T}{T_{0}}+2\pi T\sum_{\omega_{n}>0}\frac{1}{\omega_{n}}\left[1-\frac{\omega_{n}^{2}}{\omega_{n}^{2}+V^{2}}\right]=0,$ (11)
@@
are identical to each other. Here $T_{0}$ is the transition temperature at $\bm{H}=\bm{\alpha}=0$. The phase boundary between the normal and superconducting states in Fig. 1(a) is determined by two methods: solving the gap equation in Eq. (11) and finding the minimum of the free-energy in Eq. (6). Two results are identical to each other when the transition to the superconducting states is a second-order. For first-order transitions, the phase boundary is obtained from the free-energy minima. The arrow on the vertical axis indicate the Pauli limit $\mu_{\mathrm{B}}\,H_{\mathrm{P}}$ [19, 6]. The superfluid weight of the principal correlation $q_{d}(T,H_{c})$ and that for the induced odd-frequency correlation $q_{\mathrm{odd}}(T,H_{c})$ are plotted as a function of temperature in Fig. 2(a), where $H_{c}$ is obtained from the data points on the phase boundary in Fig. 1 (a) and the vertical axis is normalized to $q_{\mathrm{BCS}}$ in Eq. (S20) in SM. The transition to the superconducting state in Fig. 1(a) becomes a first-order for $T<0.556T_{0}$ [13, 22] because odd-frequency pairing correlation decreases $q(T_{c},H_{c})$ to be negative as demonstrated in Fig. 2(a). A Zeeman field seriously suppresses spin-triplet superconductivity in the parallel configuration.
Figure 2: In (a), the two superfluid weights for the parallel configuration are plotted along the phase boundary in Fig. 1(a). In (b), the three superfluid weights are plotted as a function of Zeeman fields for the perpendicular configuration, where the vertical doted line indicates the critical field at $T=0.5T_{0}$ and $\alpha=1.5T_{0}$ in Fig. 1(b).
@@
Secondly, we consider a configuration $\bm{\alpha}\perp\bm{d}$, $\bm{\alpha}\perp\bm{H}$, and $\bm{H}\perp\bm{d}$, where the three vectors are perpendicular to one another. The anomalous Green’s function near the transition temperature is supplied in Eq. (S15) in SM [3]. Here we show the results after carrying out the summation over $\bm{k}$ by shifting the wavenumber $\bm{k}\to\bm{k}\mp\bm{K}$ for $\xi_{\pm}$,
$\frac{1}{V_{\mathrm{vol}}}\sum_{\bm{k}}$ $\check{F}_{\perp}(\bm{k},\omega_{n})=\frac{N_{0}\pi}{|\omega_{n}|(\omega_{n}^{2}+\alpha^{2}+V^{2})}$
$\times\left[(\omega_{n}^{2}+V^{2})\bm{d}\cdot\hat{\bm{\sigma}}+\omega_{n}(\bm{\alpha}\times\bm{d})\cdot\hat{\bm{\sigma}}\hat{\rho}_{z}\right$.
$\left.+i(\bm{\alpha}\times\bm{d})\cdot\bm{V}\hat{\rho}_{z}\right]\hat{\sigma}_{y}\,\hat{\rho}_{y}.$ (12)
@@
The first term is the principal pairing correlation. The odd-frequency pairing correlation at the second term is induced by the SOI and suppresses superconductivity [8]. We will show that the third term listed at the bottom row in Table 1 plays a key role in the ZFIS. The linearized gap equation results in
$\ln\frac{T}{T_{0}}$ $+2\pi T\sum_{\omega_{n}>0}\frac{1}{\omega_{n}}\left[1-\frac{\omega_{n}^{2}+V^{2}}{\omega_{n}^{2}+V^{2}+\alpha^{2}}\right]=0.$ (13)
@@
Equation (13) suggests that the large SOI $\alpha>\mu_{\mathrm{B}}H_{\mathrm{P}}$ deletes superconductivity at $\bm{H}=0$. In the HTH-T phase diagram in Fig. 1(b), superconductivity is absent at $\bm{H}=0$ because we choose such large SOI as $\alpha=1.5T_{0}$ and $2.5T_{0}$. Equation (13) also indicates that a Zeeman field screens such negative effects due to the SOI. Fig. 1(b) shows that the superconducting phase appears in large Zeeman fields and that the critical temperature increases with increasing Zeeman fields. We explain the reasons below. The analytic expressions of the superfluid weight of the principal correlation $q_{d}$, that of odd-frequency pairs $q_{\mathrm{odd}}$, and that of spin-singlet pairs $q_{\perp}$ are supplied in Eqs. (S22)-(S23) in SM. They are plotted as a function of the Zeeman potential in Fig. 2(b), where we choose $\alpha=1.5T_{0}$ and fix the temperature at $T=0.5T_{0}$. The results show that $q_{\perp}$ compensates the negative weight of $q_{\mathrm{odd}}$ in Zeeman fields. Namely, the induced spin-singlet Cooper pairs stabilize the spin-triplet superconductivity in Zeeman fields. The increase of $q_{d}$ with Zeeman fields is a result of such compensation, which explains the ZFIS described by the gap equation in Eq. (13). We conclude that the interplay between the SOI and the Zeeman field realizes ZFIS. Eq. (13) suggests that $T_{c}$ goes to the $T_{0}$ for $V\gg\alpha$. However, our phenomenological theory cannot predict such limiting behavior. We have assumed in high magnetic fields that the attractive interaction between two electrons remains unchanged and that other conduction bands do not come to the Fermi level.
@@
Finally, we consider a mixed situation between the parallel and the perpendicular configurations,
$\bm{\alpha}=$ $\alpha_{x}\bm{e}_{x}+\alpha_{y}\bm{e}_{y},\quad\bm{d}=d_{x}\bm{e}_{x}+d_{z}\bm{e}_{z},$ (14)